The given equation involves a power of the variable. Find all real solutions of the equation. (If there is no real solution, enter NO REAL SOLUTION.)
x^6/5 − 729 =0

Answer by lwsshak3(11628)

The given equation involves a power of the variable. Find all real solutions of the equation. (If there is no real solution, enter NO REAL SOLUTION.)
x^6/5 − 729 =0
x^6/5=729
raise both sides to the 5/6 power
x=729^5/6=6th root of 729 raised to the 5th power
=3^5
=243